The current price of a non-dividend-paying stock is $30. Over the next six months it is expected to rise to $36 or fall to $26. Assume that the risk-free rate is 10%. What, to the nearest cent, is the value of a 6-month European call option on the stock with a strike price of $33?
Answer #1
Upmove (U)= High price/current price=36/30=1.2
Down move (D)= Low price/current price=26/30=0.8667
Risk neutral probability for up move
q = (e^(risk free rate*time)-D)/(U-D)
=(e^(0.1*0.5)-0.8667)/(1.2-0.8667)=0.55381
Call option payoff at high price (payoff H)
=Max(High price-strike price,0)
=Max(36-33,0)
=Max(3,0)
=3
Call option payoff at low price (Payoff L)
=Max(Low price-strike price,0)
=Max(26-33,0)
=Max(-7,0)
=0
Price of call option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.1*0.5)*(0.553813*3+(1-0.553813)*0)
=1.58